# Stability of $f(Q,B)$ gravity via Dynamical System Approach: a Comprehensive Bayesian Statistical Analysis

### v4

@@ II.2 Power law $f(Q,B)$
To determine the theoretical values of the Hubble rate, we can numerically solve Eq. (22). Assuming matter behaves as a pressureless perfect fluid ($p_{\text{m}}=0$), the matter density can be expressed as $\rho_{\text{m}}=3H_{0}^{2}\Omega_{\text{m}0}(1+z)^{3}$, where $z$ denotes the cosmological redshift and $\Omega_{\text{m}0}$ is the current matter density parameter. Consequently, for the specific model under consideration, the first Friedmann equation can be written as follows:
$H^{\prime\prime}(z)=\frac{-1}{f_{0}\,n(n-1)(1+z)^{2}H(z)^{3}}\times$ (29)
[−$\left[-9f_{0}(n-2m-1)H(z)^{4}+f_{0}\,n(n-1)(1+z)\right.$
$H(z)^{3}H^{\prime}(z)-6f_{0}\left((n-1)^{2}+(2+n)m\right)\times$
$(1+z)H(z)^{3}H^{\prime}(z)+f_{0}\,n(n-1)(1+z)^{2}\times$
$H(z)^{2}H^{\prime}(z)^{2}+f_{0}(1-n+2m(1+n))(1+z)^{2}$
$H(z)^{2}H^{\prime}(z)^{2}-6^{1-n-m}H_{0}^{2}(1+z)^{3}\Omega_{\text{m}0}\left(H(z)^{2}\right)^{-m}$
@@
Equation (29) is a second-order differential equation for the function $H(z)$. To solve this equation, we need to apply suitable boundary conditions. The first boundary condition is straightforward: $H(0)=H_{0}$, which sets the present value of the Hubble parameter. To satisfy the second boundary condition, it is essential to confirm that the current rate of change of the Hubble parameter aligns with the projections of the standard $\Lambda$CDM model. This model describes the expansion of the Universe and provides a specific expansion law that the derivative should follow. By aligning the first derivative of $H(z)$ with this expansion law, we can accurately determine the second initial condition needed to solve the differential equation.
@@
After taking the derivative of the equation with respect to $z$, we can derive the second initial condition for Eq. (29) as $H^{\prime}(0)=\frac{3}{2}H_{0}\Omega_{\text{m}0}$.
III Observational data, methodology and constraints
To model the Universe accurately, we require robust observational data and effective parameter estimation methodologies. Within this framework, we detail the observational datasets and methods used to constrain the model parameters $f_{0}$, $m$, and $n$. Our analysis includes a comprehensive array of data, such as CC, Pantheon+ and BAO from SNe Ia observations. By leveraging these diverse datasets, we effectively narrow down the model parameters, facilitating an in-depth exploration of the evolution of the Universe. Additionally, we explore $f(Q,B)$ gravity and its solutions involving the Hubble parameter. The CC dataset, known for its reliability and model independence, provides Hubble parameters by measuring the age difference between two passively evolving galaxies. This method allows us to derive the Hubble function at various redshifts up to $z\approx 2$. The shape of $H(z)$ is further constrained by multiple sources, including 32 data points from Hubble datasets, BAO data from various sources, and CMB data from Planck 2018. The employed methodology, utilized data, and outcomes are detailed in subsequent sections.
III.1 Cosmic Chronometers (CC)
@@
where, $\dot{a}$ is the derivative of the scale factor $a$ with respect to time $t$, and $\Delta z$ and $\Delta t$ are the differences in redshift and time, respectively, between two objects. The value of $\Delta z$ can be determined via a spectroscopic survey, while the differential ages $\Delta t$ of passively evolving galaxies can be used to estimate the value of $H(z)$. Compiling such observations can be regarded as a CC sample. We use 32 objects covering the redshift range $0.07\leq z\leq 1.965$ [39, 53]. For these measurements, one can construct a $\chi^{2}_{\text{CC}}$ estimator as follows:
@@
where, $H_{\text{obs}}$ and $H_{\text{th}}$ represent the observational and theoretical Hubble parameter values at redshift $z_{i}$, respectively, with $\sigma_{H}$ being the error in the observational value.
Figure 1: The contour plots display the $1-\sigma$ and $2-\sigma$ uncertainty regions for the model parameters $H_{0}$, $\Omega_{\text{m}0}$, $f_{0}$, $m$ and $n$. These contours are based on the combined CC + Pantheon+ datasets.
Figure 2: The contour plots display the $1-\sigma$ and $2-\sigma$ uncertainty regions for the model parameters $H_{0}$, $\Omega_{\text{m}0}$, $f_{0}$, $m$ and $n$. These contours are based on the combined CC + Pantheon+ + BAO datasets.
III.2 Supernovae type Ia (SNe Ia)
We will also consider the Pantheon+ SNe Ia data compilation, consisting of 1701 SNe Ia relative luminosity distance measurements spanning the redshift range of $2.26130.00122<z<$ 2.2613 [54]. The Pantheon+ dataset contains distance moduli estimated from 1701 light curves of 1550 spectroscopically confirmed SNe Ia with a redshift range, acquired from 18 distinct surveys. Notably, 77 of the 1701 light curves are associated with Cepheid-containing galaxies. The Pantheon+ dataset has the benefit of being able to constrain $H_{0}$ in addition to the model parameters. To fit the parameter of the model from the Pantheon+ samples, we extremie the $\chi^{2}$ function. To calculate the chi-square $(\chi^{2}_{\text{SNe}})$ value, which helps estimate the best-fit parameters using supernovae from the Pantheon compilation consisting of 1701 data points, we use the following formula:
@@ III.2 Supernovae type Ia (SNe Ia)
Specifically, the distance modulus is defined as the difference between an observed apparent magnitude of object, $m$, and its absolute magnitude, $M$, which measures its intrinsic brightness. At redshift $z_{i}$, the distance modulus is expressed as follows:
@@
where $E(z)=\frac{H(z)}{H_{0}}$ and $c$ stand for the speed of light. Additionally, the residual distance is denoted by
@@
where $\mu_{k}^{cd}$ is the Cepheid host-galaxy distance that SH0ES revealed. This covariance matrix can be coupled with the SNe covariance matrix to construct the covariance matrix for the Cepheid host galaxy. Equipped with statistical and systematic uncertainties from the Pantheon+ dataset, the combined covariance matrix is represented as $C^{\text{SNe}}_{\text{Sys}+\text{Stat}}+C^{cd}_{\text{Sys}+\text{Stat}}$. The expression above defines the $\chi^{2}$ function for the combined covariance matrix used to constrain cosmological models in the analysis.
@@
Figure 3: In the upper panel, the black error bars represent the uncertainty associated with the 32 data points from the CC sample. The solid teal line corresponds to the model, while the dashed red line represents the $\Lambda$CDM. Moving to the lower panel, we observe a red line that depicts the plot of the distance modulus of model $\mu(z)$ against redshift $z$. This teal line demonstrates a superior fit to the 1701 data points from the Pantheon+ dataset, including their associated error bars.
@@
III.3 Baryon Acoustic Oscillation (BAO)
Various surveys have significantly advanced the analysis of cosmic distance scales through BAO, including the 6-degree Field Galaxy Survey, Sloan Digital Sky Survey (SDSS), and WiggleZ Dark Energy Survey. These surveys have measured BAO signals at multiple redshifts, providing critical data for understanding the expansion of the Universe. The horizon of sound during the radiation drag era, denoted as $r_{s}(z_{*})$, is calculated using the following equation:
@@ III.3 Baryon Acoustic Oscillation (BAO)
where $c$ represents the speed of light, and $\Omega_{b0}$ and $\Omega_{\gamma 0}$ are the present baryon and photon densities, respectively.
To derive BAO constraints, the angular diameter distance $D_{A}(z)$, dilation scale $D_{v}(z)$ and the Hubble parameter $H(z)$ are used, defined by the equations:
@@
For the MCMC analysis, the same walkers, steps and priors as in the CC sample are utilized. The chi-square function for BAO is expressed as:
@@
where $X$ and $C^{-1}$ are the data vector and the inverse covariance matrix, respectively defined in Ref. [55].
Figure 5: Behavior of the deceleration parameter (upper panel) and Equation of State (EoS) parameter (lower panel) using the CC + Pantheon+ and CC + Pantheon+ + BAO datasets, with the mean values of parameters $f_{0}$, $m$, and $n$ as listed in Table 1.
@@
Figs. 1 and 2 provide the contour plots with $1-\sigma$ and $2-\sigma$ errors for the CC + Pantheon+ and CC + Pantheon+ + BAO data sets, respectively. In the top panel of Fig. 3, the evolution of the Hubble parameter as a function of redshift is illustrated. This figure compares the predictions of two models: the $\Lambda$CDM model and the $H(z)$ model derived from the numerical approach proposed in this study (depicted by the teal line), alongside observational data. One line is included for the $\Lambda$CDM model to facilitate the comparison. Additionally, the dashed-red line (labeled as $\Lambda$CDM) is derived from the standard prediction of the $\Lambda$CDM model with the parameters $H_{0}$, and $\Omega_{\text{m}}$. In Fig. 3, the bottom panel shows a comparison of the distance modulus using our $f(Q,B)$ model (teal line) and the $\Lambda$CDM model (dashed red line) predictions. Both models were considered with their respective parameters. The similarity between our model and $\Lambda$CDM prediction is evident. However, when the same shared parameter values were used, the models deviated from each other, mainly in the apparent magnitude prediction. The red dashed line in Fig. 3 shows that our model fits better with the $\Lambda$CDM. Also, the model accurately captures the behavior of the Hubble function, as shown by the consistency of error bars. In Fig. 3, the observed distance modulus of the 1701 SNe Ia dataset is depicted, along with the best-fit theoretical curves of the distance modulus function $\mu(z)$ shown as a teal line.
@@
Fig. 4 displays the best-fit values and associated uncertainties for the model parameters $H_{0}$, $\Omega_{\text{m}0}$, $f_{0}$, $m$, and $n$, derived from our MCMC analysis. The plot visually represents the parameter ranges obtained from different datasets, including CC + Pantheon+ and CC + Pantheon+ + BAO. The Hubble constant $H_{0}$ values range from approximately 68.913 $\text{Km}\,\,\text{s}^{-1}\,\text{Mpc}^{-1}$ to 72.426 $\text{Km}\,\,\text{s}^{-1}\,\text{Mpc}^{-1}$, while the matter density parameter $\Omega_{\text{m}0}$ spans from 0.248 to 0.312. The parameters $f_{0}$, $m$, and $n$ exhibit ranges of 1.078 to 1.475, −$-2.127$ to −$-1.89$, and 2.876 to 3.144, respectively. These ranges highlight the variability and discrepancies in the parameter estimates, underscoring the robustness and reliability of the model fits to the observational data. The whisker plot effectively conveys the uncertainties inherent in the model parameters, providing a comprehensive overview of the results from the MCMC analysis.
@@
Fig. 5 illustrates the significance of the deceleration parameter $q$, a crucial metric in cosmology that provides insights into the dynamics of the Universe. A positive $q$ indicates deceleration, while a negative $q$ signifies acceleration. Analysis of the CC + Pantheon+ and CC + Pantheon+ + BAO datasets reveals that $q$ transitions from positive in the past, indicating early deceleration, to negative in the present, indicating current acceleration, as depicted in Fig. 5. At the current cosmic epoch, the deceleration parameter $q_{0}$ has been measured as −$-0.506$ and −$-0.549$ for the CC + Pantheon+ and CC + Pantheon+ + BAO datasets, respectively. These values are in good agreement with the range of $q_{0}=-0.528^{+0.092}_{-0.088}$ determined by recent observations [56]. Current observations align with this deceleration parameter, and the derived model demonstrates a smooth transition from deceleration to acceleration at $z_{t}=0.763$ and $z_{t}=0.67$ for the CC + Pantheon+ and CC + Pantheon+ + BAO datasets, respectively. The recovered transition redshift $z_{t}$ is consistent with current constraints based on 11 $H(z)$ observations reported by Busca et al. [57] for redshifts $0.2\leq z\leq 2.3$, $z_{t}=0.74\pm 0.5$ from Farooq et al. [58], $z_{t}=0.7679^{+0.1831}_{-0.1829}$ by Capozziello et al. [59], and $z_{t}=0.60^{+0.21}_{-0.12}$ by Yang et al. [60]. Similarly, the EoS parameter $(\omega_{\text{DE}})$, is integral to understanding the evolution of the Universe, as it correlates with the energy sources influencing this progression. The current EoS values for DE, represented by $\omega_{\text{DE}}(z=0)$, are determined to be −$-1.032$ and −$-1.004$ for the CC + Pantheon+ and CC + Pantheon+ + BAO datasets, respectively. Various cosmological studies have also placed constraints on the EoS parameter. For instance, the Planck 2018 results yielded $\omega_{\text{DE}}=-1.03\pm 0.03$ [61], and the WAMP + CMB analysis reported $\omega_{\text{DE}}=-1.079^{+0.090}_{-0.089}$ [62]. By computing the associated energy density and pressure of DE, we can observe the fluctuations in the effective DE EoS, which are depicted in redshift [Fig. 5].
@@
We evaluate the models against the standard $\Lambda$CDM model using the Akaike Information Criterion (AIC) and the Bayesian Information Criterion (BIC), in addition to $\chi^{2}_{\text{min}}$. Both AIC and BIC consider the model’s goodness of fit and its complexity, which depends on the number of parameters $(n)$. The AIC is calculated as:
@@
In statistical modeling, a lower AIC value indicates a better fit to the data, accounting for model complexity. This penalizes models with more parameters, even if they fit the data better. The BIC is computed as:
@@
[…]
@@
IV Dynamical System Analysis
The methods of dynamical systems are valuable for analyzing the overall long-term dynamics of a particular cosmological model. This involves an equation, $x^{\prime}=f(x)$, where $x$ is a column vector and $f(x)$ is the equivalent vector of autonomous equations. In this method, the prime symbol represents the derivative with respect to the number of e-folding, $N=\text{ln}\,a(t)$. The general form of the dynamical system for the modified FLRW equations defined by Eq. (22) can be generated through this approach. Let us define a new variable:
$X=f_{B},\hskip 8.5359ptY=\frac{\dot{f}_{B}}{H},\hskip 8.5359ptZ=\frac{\dot{H}}{H^{2}},\hskip 8.5359ptV=\frac{\kappa\,\rho_{\text{r}}}{3H^{2}},$
$W=-\frac{f}{6H^{2}},\hskip 8.5359pt\Omega_{\text{m}}=\frac{\kappa\,\rho_{\text{m}}}{3H^{2}},\hskip 8.5359pt\Omega_{\text{DE}}=\frac{\kappa\,\rho_{\text{DE}}}{3H^{2}}\,.$ (45)
Thus, from Eq. (22), we have the algebraic identity
[…]
together with the density parameters:

### v5

@@ 2.2 Power law $f(Q,B)$
To determine the theoretical values of the Hubble rate, we can numerically solve Equation (22). Assuming matter behaves as a pressureless perfect fluid ($p_{\text{m}}=0$), the matter density can be expressed as $\rho_{\text{m}}=3H_{0}^{2}\Omega_{\text{m}0}(1+z)^{3}$, where $z$ denotes the cosmological redshift; it is defined as $\frac{a}{a_{0}}=\frac{1}{1+z}$, where $a_{0}$ is the scale factor at present and $\Omega_{\text{m}0}$ is the current matter density parameter. Consequently, for the specific model under consideration, the first Friedmann equation can be written as follows:
$H^{\prime\prime}(z)$ $=$ $\frac{-1}{f_{0}\,n(n-1)(1+z)^{2}H(z)^{3}}\left[-9f_{0}(n-2m-1)H(z)^{4}+f_{0}\,n(n-1)(1+z)\right.H(z)^{3}H^{\prime}(z)$ (29)
$-6f_{0}\left((n-1)^{2}+(2+n)m\right)(1+z)H(z)^{3}H^{\prime}(z)+f_{0}\,n(n-1)(1+z)^{2}H(z)^{2}H^{\prime}(z)^{2}$
$+f_{0}(1-n+2m(1+n))(1+z)^{2}H(z)^{2}H^{\prime}(z)^{2}-6^{1-n-m}H_{0}^{2}(1+z)^{3}\Omega_{\text{m}0}\left(H(z)^{2}\right)^{-m}$
@@
In this study, we examine a second-order nonlinear ordinary differential equation that presents significant challenges for analytical solutions. As a result, we depend to numerical methods to solve Equation (29) in order to obtain theoretical values for the Hubble rate. To solve this equation, we need to apply suitable boundary conditions. The first boundary condition is straightforward: $H(0)=H_{0}$, which sets the present value of the Hubble parameter. To satisfy the second boundary condition, it is essential to confirm that the current rate of change of the Hubble parameter aligns with the projections of the standard $\Lambda$CDM model. This model describes the expansion of the Universe and provides a specific expansion law that the derivative should follow. By aligning the first derivative of $H(z)$ with this expansion law, we can accurately determine the second initial condition needed to solve the differential equation:
@@
After taking the derivative of the Equation (30) with respect to $z$, we can derive the second initial condition for Equation (29) as $H^{\prime}(0)=\frac{3}{2}H_{0}\Omega_{\text{m}0}$.
3 Observational data, methodology and constraints
To model the Universe accurately, we require robust observational data and effective parameter estimation methodologies. Within this framework, we detail the observational data sets and methods used to constrain the model parameters $f_{0}$, $m$, and $n$. Our analysis includes a comprehensive array of data, such as CC, Pantheon+ and BAO from SNe Ia observations. We conducted MCMC analyses by integrating results from the CC, Pantheon+, and BAO data sets, assuming a correlation among them. Our findings are presented for the CC + Pantheon+ and CC + Pantheon+ + BAO combinations, thereby reflecting their interdependence. By leveraging these diverse data sets, we effectively narrow down the model parameters, facilitating an in-depth exploration of the evolution of the Universe. Additionally, we explore $f(Q,B)$ gravity and its solutions involving the Hubble parameter. The CC data set, known for its reliability and model independence, provides Hubble parameters by measuring the age difference between two passively evolving galaxies. This method allows us to derive the Hubble function at various redshifts up to $z\approx 2$. The shape of $H(z)$ is further constrained by multiple sources, including 32 data points from Hubble data sets, BAO data from various sources, and CMB data from Planck 2018. The employed methodology, utilized data, and outcomes are detailed in subsequent sections.
3.1 Cosmic Chronometers (CCs)
@@
where $\dot{a}$ is the derivative of the scale factor $a$ with respect to time $t$, and $\Delta z$ and $\Delta t$ are the differences in redshift and time, respectively, between two objects. The value of $\Delta z$ can be determined via a spectroscopic survey, while the differential ages $\Delta t$ of passively evolving galaxies can be used to estimate the value of $H(z)$. Compiling such observations can be regarded as a CC sample. We use 32 objects covering the redshift range $0.07\leq z\leq 1.965$ (Moresco et al., 2022; Lohakare et al., 2023b). For these measurements, one can construct a $\chi^{2}_{\text{CC}}$ estimator as follows:
@@
where $H_{\text{obs}}$ and $H_{\text{th}}$ represent the observational and theoretical Hubble parameter values at redshift $z_{i}$, respectively, with $\sigma_{H}$ being the error in the observational value.
3.2 Supernovae type Ia (SNe Ia)
We will also consider the Pantheon+ SNe Ia data compilation, consisting of 1701 SNe Ia relative luminosity distance measurements spanning the redshift range of $2.26130.00122<z<$ 2.2613 (Brout et al., 2022). The Pantheon+ data set contains distance moduli estimated from 1701 light curves of 1550 spectroscopically confirmed SNe Ia with a redshift range acquired from 18 distinct surveys. Notably, 77 of the 1701 light curves are associated with Cepheid-containing galaxies. The Pantheon+ data set has the benefit of being able to constrain $H_{0}$ in addition to the model parameters. To fit the parameter of the model from the Pantheon+ samples, we minimize the $\chi^{2}$ function. To calculate the chi-square $(\chi^{2}_{\text{SNe}})$ value, which helps estimate the best-fit parameters using SNe from the Pantheon compilation consisting of 1701 data points, we use the following formula:
@@ 3.2 Supernovae type Ia (SNe Ia)
Figure 1: The contour plots display the $1\sigma$ and $2\sigma$ uncertainty regions for the model parameters $H_{0}$, $\Omega_{\text{m}0}$, $f_{0}$, $m$ and $n$. These contours are based on the combined CC + Pantheon+ data sets.
Figure 2: The contour plots display the $1\sigma$ and $2\sigma$ uncertainty regions for the model parameters $H_{0}$, $\Omega_{\text{m}0}$, $f_{0}$, $m$ and $n$. These contours are based on the combined CC + Pantheon+ + BAO data sets.
Specifically, the distance modulus is defined as the difference between an observed apparent magnitude of an object, $m$, and its absolute magnitude, $M$, which measures its intrinsic brightness. At redshift $z_{i}$, the distance modulus is expressed as follows:
@@
where $E(z)=\frac{H(z)}{H_{0}}$ and $c$ stands for the speed of light. Additionally, the residual distance is denoted by
@@
where $\mu_{k}^{cd}$ is the Cepheid host-galaxy distance that SH0ES revealed. This covariance matrix can be coupled with the SNe covariance matrix to construct the covariance matrix for the Cepheid host galaxy. Equipped with statistical and systematic uncertainties from the Pantheon+ data set, the combined covariance matrix is represented as $C^{\text{SNe}}_{\text{Sys}+\text{Stat}}+C^{cd}_{\text{Sys}+\text{Stat}}$. The expression above defines the $\chi^{2}$ function for the combined covariance matrix used to constrain cosmological models in the analysis:
@@
Figure 3: In the left panel, the black error bars represent the uncertainty associated with the 32 data points from the CC sample. The solid teal line corresponds to the model, while the dashed red line represents the $\Lambda$CDM. Moving to the right panel, we observe a red line that depicts the plot of the distance modulus of model $\mu(z)$ against redshift $z$. The teal line demonstrates a superior fit to the 1701 data points from the Pantheon+ data set, including their associated error bars.
@@
3.3 Baryon Acoustic Oscillation (BAO)
Various surveys have significantly advanced the analysis of cosmic distance scales through BAO, including the 6-degree Field Galaxy Survey, Sloan Digital Sky Survey, and WiggleZ Dark Energy Survey. These surveys have measured BAO signals at multiple redshifts, providing critical data for understanding the expansion of the Universe. The horizon of sound during the radiation drag era, denoted as $r_{s}(z_{*})$, is calculated using the following equation:
@@ 3.3 Baryon Acoustic Oscillation (BAO)
where $c$ represents the speed of light and $\Omega_{b0}$ and $\Omega_{\gamma 0}$ are the present baryon and photon densities, respectively.
To derive BAO constraints, the angular diameter distance $D_{A}(z)$, dilation scale $D_{v}(z)$ and the Hubble parameter $H(z)$ are used, defined by the equations
@@
For the MCMC analysis, the same walkers, steps, and priors as in the CC sample are utilized. The chi-square function for BAO is expressed as
@@
where $X$ and $C^{-1}$ are the data vector and the inverse covariance matrix, respectively defined in Ref. (Giostri et al., 2012).
Figure 5: Behavior of the deceleration parameter (left panel) and EoS parameter (right panel) using the CC + Pantheon+ and CC + Pantheon+ + BAO data sets, with the mean values of parameters $f_{0}$, $m$, and $n$ as listed in Table 1.
@@
Figures 1 and 2 provide the contour plots with $1\sigma$ and $2\sigma$ errors for the CC + Pantheon+ and CC + Pantheon+ + BAO data sets, respectively. The values obtained from the MCMC analysis that best fit the model are shown in Table 1. In the left panel of Figure 3, the evolution of the Hubble parameter as a function of redshift is illustrated. This figure compares the predictions of two models: the $\Lambda$CDM model and the $H(z)$ model derived from the numerical approach proposed in this study (depicted by the teal line), alongside observational data. One line is included for the $\Lambda$CDM model to facilitate the comparison. Additionally, the dashed red line (labeled as $\Lambda$CDM) is derived from the standard prediction of the $\Lambda$CDM model with the parameters $H_{0}$, and $\Omega_{\text{m}}$. In Figure 3, the right panel shows a comparison of the distance modulus using our $f(Q,B)$ model (teal line) and the $\Lambda$CDM model (dashed red line) predictions. Both models were considered with their respective parameters. The similarity between our model and the $\Lambda$CDM prediction is evident. However, when the same shared parameter values were used, the models deviated from each other, mainly in the apparent magnitude prediction. The dashed red line in Figure 3 shows that our model fits better with the $\Lambda$CDM. Also, the model accurately captures the behavior of the Hubble function, as shown by the consistency of the error bars. In Figure 3, the observed distance modulus of the 1701 SNe Ia data set is depicted, along with the best-fit theoretical curves of the distance modulus function $\mu(z)$ shown as a teal line.
@@
Figure 4 displays the best-fit values and associated uncertainties for the model parameters $H_{0}$, $\Omega_{\text{m}0}$, $f_{0}$, $m$, and $n$, derived from our MCMC analysis. The plot visually represents the parameter ranges obtained from different data sets, including CC + Pantheon+ and CC + Pantheon+ + BAO. The Hubble constant $H_{0}$ values range from approximately 68.913 $\text{Km}\,\,\text{s}^{-1}\,\text{Mpc}^{-1}$ to 72.426 $\text{Km}\,\,\text{s}^{-1}\,\text{Mpc}^{-1}$, while the matter density parameter $\Omega_{\text{m}0}$ spans from 0.248 to 0.312. The parameters $f_{0}$, $m$, and $n$ exhibit ranges of 1.078 to 1.475, $-2.127$ to $-1.89$, and 2.876 to 3.144, respectively. These ranges highlight the variability and discrepancies in the parameter estimates, underscoring the robustness and reliability of the model fits to the observational data. The whisker plot effectively conveys the uncertainties inherent in the model parameters, providing a comprehensive overview of the results from the MCMC analysis. This comparative analysis not only enhances our understanding of the data but also facilitates informed decision-making based on observed patterns.
@@
Figure 5 illustrates the significance of the deceleration parameter $q$, a crucial metric in cosmology that provides insights into the dynamics of the Universe. A positive $q$ indicates deceleration, while a negative $q$ signifies acceleration. Analysis of the CC + Pantheon+ and CC + Pantheon+ + BAO data sets reveals that $q$ transitions from positive in the past, indicating early deceleration, to negative in the present, indicating current acceleration, as depicted in Figure 5. At the current cosmic epoch, the deceleration parameter $q_{0}$ has been measured as $-0.506$ and $-0.549$ for the CC + Pantheon+ and CC + Pantheon+ + BAO data sets, respectively. These values are in good agreement with the range of $q_{0}=-0.528^{+0.092}_{-0.088}$ determined by recent observations (Gruber & Luongo, 2014). Current observations align with this deceleration parameter, and the derived model demonstrates a smooth transition from deceleration to acceleration at $z_{t}=0.763$ and $z_{t}=0.67$ for the CC + Pantheon+ and CC + Pantheon+ + BAO data sets, respectively. The recovered transition redshift $z_{t}$ is consistent with current constraints based on 11 $H(z)$ observations reported by Busca et al. (2013) for redshifts $0.2\leq z\leq 2.3$, $z_{t}=0.74\pm 0.5$ from Farooq & Ratra (2013), $z_{t}=0.7679^{+0.1831}_{-0.1829}$ by Capozziello et al. (2014), and $z_{t}=0.60^{+0.21}_{-0.12}$ by Yang & Gong (2020). Similarly, the EoS parameter $(\omega_{\text{DE}})$, is integral to understanding the evolution of the Universe, as it correlates with the energy sources influencing this progression. The current EoS values for DE, represented by $\omega_{\text{DE}}(z=0)$, are determined to be $-1.032$ and $-1.004$ for the CC + Pantheon+ and CC + Pantheon+ + BAO data sets, respectively. Various cosmological studies have also placed constraints on the EoS parameter. For instance, the Planck 2018 results yielded $\omega_{\text{DE}}=-1.03\pm 0.03$ (Aghanim et al., 2020b), and the WAMP + CMB analysis reported $\omega_{\text{DE}}=-1.079^{+0.090}_{-0.089}$ (Hinshaw et al., 2013). By computing the associated energy density and pressure of DE, we can observe the fluctuations in the effective DE EoS, which are depicted in redshift (Figure 5).
@@
We evaluate the models against the standard $\Lambda$CDM model using the Akaike Information Criterion (AIC) and the Bayesian Information Criterion (BIC), in addition to $\chi^{2}_{\text{min}}$. Both the AIC and BIC consider the model’s goodness of fit and its complexity, which depends on the number of parameters $(n)$. The AIC is calculated as
@@
In statistical modeling, a lower AIC value indicates a better fit to the data, accounting for model complexity. This penalizes models with more parameters, even if they fit the data better. The BIC is computed as
@@
where $\mathcal{N}$ is the number of data samples used in the MCMC process. The ${}_{\text{c}}$) is defined as
@@
4 Dynamical System Analysis
The methods of dynamical systems are valuable for analyzing the overall long-term dynamics of a particular cosmological model. This involves an equation, $x^{\prime}=f(x)$, where $x$ is a column vector and $f(x)$ is the equivalent vector of autonomous equations. In this method, the prime symbol represents the derivative with respect to the number of $e$–folding, $N=\text{ln}\,a(t)$. The general form of the dynamical system for the modified FLRW equations defined by Equation (22) can be generated through this approach. Let us define a new variable:
$X=f_{B},\hskip 8.5359ptY=\frac{\dot{f}_{B}}{H},\hskip 8.5359ptZ=\frac{\dot{H}}{H^{2}},\hskip 8.5359ptV=\frac{\kappa\,\rho_{\text{r}}}{3H^{2}},\hskip 8.5359ptW=-\frac{f}{6H^{2}},\hskip 8.5359pt\Omega_{\text{m}}=\frac{\kappa\,\rho_{\text{m}}}{3H^{2}},\hskip 8.5359pt\Omega_{\text{DE}}=\frac{\kappa\,\rho_{\text{DE}}}{3H^{2}}\,.$ (45)
Thus, from Equation (22), we have the algebraic identity
$\Omega_{\text{m}}+\Omega_{\text{r}}+\Omega_{\text{DE}}=1\,,$ (46)
together with the density parameters